OpenStudy (anonymous):

THIS IS SOOOO HARD!! HLP PLZ!!!!!!!!!!!!!!!! i cant do it!!! aaaaggghhh! One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

i need help:(

OpenStudy (turingtest):

let #people =x and #horses =y you need one expression for the total number of heads, and another for the total number of legs what is the equation for the total number of heads in terms of x and y ?

OpenStudy (anonymous):

uhhhh...if x=the heads, then there would be 74, and if y was the legs there would be 196...

OpenStudy (turingtest):

x doesn't equal heads, it equals number of people

OpenStudy (turingtest):

we are counting heads of both horses and humans

OpenStudy (anonymous):

oh.. then.. well there is no people yet..

OpenStudy (turingtest):

if x is the number of humans, and y the number of horses, and the number of heads total is 74, how can we write that mathematically?

OpenStudy (anonymous):

wouldnt u have 2 solve 4 x?

OpenStudy (turingtest):

eventually, but first you need to do what I'm saying: get one equation for the number of heads, and another for the number of legs

OpenStudy (anonymous):

i dont know how!:(

OpenStudy (turingtest):

there are x humans there are y horses okay?

OpenStudy (anonymous):

oh.. ok that was a lot easier than wat i was thinking

OpenStudy (anonymous):

sorry:)

OpenStudy (anonymous):

never mind, this was extra credit anyways so i don't really HAVE to do it..:/

OpenStudy (turingtest):

we have not solved the problem at all

OpenStudy (anonymous):

/popcorn

OpenStudy (anonymous):

i no but its very frusturating..

OpenStudy (turingtest):

I'll let someone else have a crack at explaining it.

OpenStudy (anonymous):

i dont HAVE 2 do it..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks 4 trying though;)

OpenStudy (anonymous):

\[\ \ \ n_1 - \text{the number of humans}\\\ \ \ n_2-\text{the number of horses}\\\text{Now, we know that both humans and horses only have one head, and we're given}\\\text{that there are 74 heads.}\\\ \ \ n_1+n_2=74\\\text{We know that humans have 2 legs,while horses have 4. There's a total of 196 legs.}\\\ \ \ 2n_1+4n_2=196\\\ \ \ n_1+2n_2=98\\\text{Now, we have to eliminate a variable first to solve.}\\\ \ \ n_1=98-2n_2\\\text{Now we can substitute in our first equation.}\\\ \ \ 98-2n_2+n_2=74\\\ \ -n_2=-24\\\ \ \ n_2=24\\\text{Since we have solved for the number of horses, we can determine the number of humans.}\\\ \ \ n_1+24=74\\\ \ \ n_1=50\]

OpenStudy (anonymous):

wow...

OpenStudy (anonymous):

so is the answer 50?